# Su-RoBERTa: A Semi-supervised Approach to Predicting Suicide Risk through Social Media using Base Language Models

### v1

@@ 2 Literature Review
In recent years, to prevent suicide, several researchers have started working on AI-based solutions for suicidal risk prediction using social media posts from several social media websites such as Twitter, Reddit, etc. Considering the work done by several researchers on suicidal risk assessment, such as Mirtaheri et al., [5] proposed an AL-BTCN model consists of LSTM (Long Short Term Memory), Bi-TCN (Bidirectional Temporal Convolutional Neural Network) and self attention layer for suicide ideation detection among users. They extracted word embeddings from BERT (Bidirectional Encoder Representations from Transformers) model on Twitter and Reddit datasets obtained from Github and Kaggle and fed into AL-BTCN model for training. They achieved F1 score of 92 percent, 92 percent and 95 percent from two Twitter datasets and one Reddit dataset respectively. Another approach is proposed by Oyewale et al. [6] uses CNN and Bi-LSTM network for suicidal risk assessment. They extracted the Word2Vec and FastText word embeddings on a Reddit dataset proposed by Aldhyani et al. [7] and used it to train their proposed architecture and achieved an F1 score of 90.30 percent and 93.56 percent using Word2Vec and FastText embeddings respectively.
Squires et. al [8] leveraged semi-supervised learning and proposed deep label bayesian smoothing method for capturing uncertainties during assessment of suicidal risk and using this technique to train CNN based network. They achieved accuracy and F1 Score of 52.333 percent and 47.77 percent on Reddit C-SSRS dataset proposed by Gaur et al. [9]. Sawhney et. al [10] considered suicidal risk prediction as ordinal Regression problem and proposed SISMO model which consists of Bi-directional Long Short Term Memory (Bi-LSTM) and attention mechanism, for this purpose. They extracted the textual embeddings from the dataset proposed by Gaur et al. [9] and used it to train their proposed architecture and achieved the F score of 73 percent. Priyamvada et. al [11] extracted textual embeddings from the posts od Twitter dataset obtained from GitHub using Word2Vec and created a Stacked CNN - 2 Layer LSTM based architecture for suicidal risk assessment and trained the model on it and achieved accuracy score of 93.92 percent.
Qorich [12] et al. proposed Bi-LSTM and CNN based network which leverage the three types of embeddings that are Word2Vec, FastText and GloVe as well as they used Large Language Models (LLMs) such as GPT-2, BERT, etc. They trained their models on Reddit dataset obtained from Kaggle and achieved the F1 scores of 94.95 percent and 97.69 percent on Bi-LSTM and CNN based network and GPT-2 model respectively. Allen et al. [13] leveraged the use of Linguistic Inquiry and Word Count (LIWC) features for suicide risk prediction and and trained their proposed CNN architecture using 10 Fold Cross Validation technique and achieved macro F1 score of 50 percent on test set of dataset proposed by Shing et al. [14] during CLPsych 2019 competition [15]. Lastly, Bitew et al. [16] used weighted TD-IDF features, extracted emotion features from the text using DeepMoji model and used them to train logistic regression and SVM model and ensemble of all the features and models. They achieved the macro F1 score of 44.5 percent on test set of suicidal risk assessment dataset proposed by Shing et al. [14] during CLPsych 2019 competition [15] by training logistic regression model on weighted TF-IDF features.
@@
Interestingly, LLMs are also being applied in various other domains, such as the scientific and educational fields. In the scientific domain, LLMs have shown promise in tasks like citation generation [28, 39] and grammatical error correction [30, 31, 33], helping improve the accuracy and fluency of scientific writing. In the educational domain, their use spans across subjects like mathematics[34, 35, 36, 37] and physics[29, 32], where researchers have developed methodologies to enhance learning experiences as well as Student engagement assessment [38, 40], along with works on code generation [41] effectively bridging LLM’s capabilities with domain-specific needs.
@@
However, there are few more researches based on highlighting and summarizing the evidence of suicidal risk in social media posts such as works of Zhu et al. [17] who leveraged the use of Xiahai model, finetuned on ChatGLM3-6B model for evidence extraction task on suicide related posts on Reddit. For better evidence extraction, they created their custom prompts and leveraged the use of Chain of Thoughts Prompting, One-Shot learning, text matching and regular expressions. From the extracted evidence, they created the summaries. They used their methodology on the Reddit dataset proposed by Shing et al. [14] and Zirikly et al. [15] and achieved the good performance in all the evaluation metrics proposed in the CLPsych 2024 competition [18]. Finally, Uluslu et al. [19] leverages the use of 4-bit quantized Mistral 7B-Instruct-v0.2 model and performed prompt engineering in an iterative way to extract evidences and perform relevant string matching for extraction of correct evidence from suicidal posts of Reddit data proposed by proposed by Shing et al. [14] and Zirikly et al. [15]. They also performed summarization of evidence, they predicted the emotions of the posts using RoBERTa based emotion recognition model and summarization is performed using Prompt engineering and methodology similar to Retrieval Augmented Generation (RAG). They achieved the better performance in all the evaluation metrics proposed in the CLPsych 2024 competition [18].
@@ 3 Dataset Description
The Dataset used for the IEEE BigData 2024 Detection of Suicide Risk on Social Media Competition was proposed by Li et al. [20] and it consists of 2100 Reddit posts, out of which 2000 are reserved for model training and validation and 100 are reserved for testing purposes. The training set consists of 500 labeled Reddit posts and 1500 unlabeled Reddit posts, and the test set contains 100 posts whose labels are unknown to us as per the competition guidelines. The leaderboard position was determined by the evaluation of the revealed test set without labels, and the Final evaluation was done on an unrevealed test set.

### v2

@@ 2 Literature Review
Recently, several researchers have been working on automatic tools for the detection of suicide risk using machine learning, deep learning and base language models on the textual data obtained from social media websites. For instance, Matero et al. [5] leveraged the use of message, post ans use level statistics, BERT embeddings etc. and trained logistic regression and LSTM model with attention mechanism for suicide risk assessment. This model is also known as ”Context-BERT” in several research work (such as in [6]) and achieved F1 score of 50 per cent on the dataset proposed by Shing et al. [7] during CLPsych 2019 competition [8]. Another approach is proposed by Allen et al. [9] leveraged the use of LIWC features for predicting the level of suicidal risk and trained their proposed CNN architecture using 10 Fold Cross Validation technique and achieved a macro average F1 score of 50 percent on test set of dataset proposed by Shing et al. [7] during CLPsych 2019 competition [8].
@@
Bitew et al. [10] used weighted TD-IDF features, extracted emotion features from the text using the DeepMoji model and used them to train logistic regression and SVM model and combined all the features and models. They achieved the macro F1 score of 44.5 per cent on a test set of the suicide risk assessment dataset proposed by Shing et al. [7] during the CLPsych 2019 competition [8] by using weighted TF-IDF vectors to train the logistic regression model. Priyamvada et al. [11] extracted textual embeddings from the posts of the Twitter dataset obtained from GitHub using Word2Vec and created a series of CNN-LSTM based architecture for suicide risk assessment and trained the model on it and achieved an accuracy score of 93.92 per cent. Mirtaheri et al., [12] proposed an AL-BTCN model consisting of LSTM, Bi-TCN (Bidirectional Temporal Convolutional Neural Network) and self-attention layer for suicide ideation detection among users. They extracted word embeddings from the BERT (Bidirectional Encoder Representations from Transformers) on Twitter and Reddit datasets obtained from GitHub and Kaggle and fed them into the AL-BTCN model for training. They achieved an F1 score of 92 per cent, 92 per cent and 95 per cent from two Twitter datasets and one Reddit dataset, respectively. Finally, Sawhney et al. [6] proposed a new approach for suicide risk assessment by leveraging the use of adversarial learning to train the deep learning architecture consisting post level embeddings obtained from BERT and performed modelling using LSTMs and Bi-LSTMs and achieved the F1 score of 64 per cent on a dataset created by Gaur et al. [13].
@@
Interestingly, LLMs are also being applied in various other domains, such as the scientific and educational fields. In the scientific domain, LLMs have shown promise in tasks like citation generation [14] and grammatical error correction [15], helping improve the accuracy and fluency of scientific writing. In the educational domain, their use spans across subjects like mathematics[16] and physics[17], where researchers have developed methodologies to enhance learning experiences as well as Student engagement assessment [18].
@@
In the case of the suicide risk prediction task, there are also a few more types of research based on highlighting and summarizing the evidence of suicidal risk in social media posts, such as the works of Zhu et al. [19] who leveraged the use of Xiahai model, finetuned on ChatGLM3-6B model for evidence extraction task on suicide-related posts on Reddit. For better evidence extraction, they created their custom prompts and leveraged the use of Chain of Thoughts Prompting, One-Shot learning, text matching and regular expressions. From the extracted evidence, they created the summaries. They used their methodology on the Reddit dataset proposed by Shing et al. [7] and Zirikly et al. [8] achieving good performance in all the evaluation metrics proposed in the CLPsych 2024 competition [20]. Finally, Uluslu et al. [21] leverage the use of 4a -bit quantized Mistral 7B-Instruct-v0.2 model and perform prompt engineering in an iterative way to extract evidence and perform relevant string matching for extraction of correct evidence from suicidal posts of Reddit data proposed by Shing et al. [7] and Zirikly et al. [8]. They also performed summarization of evidence, they predicted the emotions of the posts using RoBERTa based emotion recognition model and summarization is performed using Prompt engineering and methodology similar to Retrieval Augmented Generation (RAG). They performed better in all the evaluation metrics proposed in the CLPsych 2024 competition [20].
@@ 3 Dataset Description
The Dataset used for this challenge was proposed by Li et al. [22] and it consists of 2100 Reddit posts, out of which 2000 are reserved for model training and validation and 100 are reserved for testing purposes. The training set consists of 500 labeled Reddit posts and 1500 unlabeled Reddit posts, and the test set contains 100 posts whose labels are unknown to us as per the competition guidelines. The leaderboard position was determined by the evaluation of the revealed test set without labels, and the Final evaluation was done on an unrevealed test set.